If animal is 50ft tall with length of 120ft long how heavy will it be with four legs and how big will a wingspan needed? Also how big will the wingspan be for this creature needed to be to carry its weight to fly high in sky?
1
Expert's answer
2020-06-17T09:35:26-0400

Calculate the mass of the creature assuming standard creature width a=10 ft. The standard density of flesh is 1100 kg/m3, or 67 lb/ft3.


"m=\rho V=\rho abc."

Assume that the lifting pressure is p=2000 lbf/ft2 and average wing length l=10 ft.

The force of gravity must be balanced by the lifting force:


"mg=pA,\\
mg=pls,"

the wingspan needed will be


"s=\frac{mg}{pl}=\frac{\rho abc}{pl},\\\space\\
s=\frac{67 \cdot10\cdot50\cdot120}{2000\cdot10}=201\text{ ft}."